# CTPD: Cross-Modal Temporal Pattern Discovery for Enhanced Multimodal Electronic Health Records Analysis

### v2

@@ Abstract
Integrating multimodal Electronic Health Records (EHR) data—such as numerical time series and free-text clinical reports—has great potential in predicting clinical outcomes. However, prior work has primarily focused on capturing temporal interactions within individual samples and fusing multimodal information, overlooking critical temporal patterns across patients. These patterns, such as trends in vital signs like abnormal heart rate or blood pressure, can indicate deteriorating health or an impending critical event. Similarly, clinical notes often contain textual descriptions that reflect these patterns. Identifying corresponding temporal patterns across different modalities is crucial for improving the accuracy of clinical outcome predictions, yet it remains a challenging task. To address this gap, we introduce a Cross-Modal Temporal Pattern Discovery (CTPD) framework, designed to efficiently extract meaningful cross-modal temporal patterns from multimodal EHR data. Our approach introduces shared initial temporal pattern representations which are refined using slot attention to generate temporal semantic embeddings. To ensure rich cross-modal temporal semantics in the learned patterns, we introduce a Temporal Pattern Noise Contrastive Estimation (TP-NCE) loss for cross-modal alignment, along with two reconstruction losses to retain core information of each modality. Evaluations on two clinically critical tasks—48-hour in-hospital mortality and 24-hour phenotype classification—using the MIMIC-III database demonstrate the superiority of our method over existing approaches. The code is anonymously available at https://github.com/HKU-MedAI/CTPD.
@@
Fuying Wang1∗, Feng Wu1∗, Yihan Tang1, Lequan Yu1 1Department of Statistics and Actuarial Science, School of Computing and Data Science, The University of Hong Kong {fuyingw,fengwu96,hank-tang}@connect.hku.hk, lqyu@hku.hk Correspondence: lqyu@hku.hk
@@ 1 Introduction
The increasing availability of Electronic Health Records (EHR) presents significant opportunities for advancing predictive modeling in healthcare (Acosta et al., 2022; Wang et al., 2024). EHR data is inherently multimodal and time-aware, encompassing structured data like vital signs, laboratory results, and medications, as well as unstructured data such as free-text clinical reports Kim et al. (2023). Integrating these diverse data types is crucial for comprehensive patient monitoring and accurate prediction of clinical outcomes Hayat et al. (2022); Wang et al. (2023); Zhang et al. (2023b). However, the irregularity and heterogeneity of multimodal data present significant challenges for precise outcome prediction.
@@
Existing approaches primarily address either the irregularity of data Shukla and Marlin (2019); Horn et al. (2020); Zhang et al. (2021, 2023b) or the fusion of multiple modalities Huang et al. (2020); Zhang et al. (2020b); Xu et al. (2021); Kline et al. (2022), but they often neglect broader temporal trends that span across patient cases. These cross-modal temporal patterns, present in both structured and unstructured data, can provide high-level semantic insights into a patient’s health trajectory and potential risks Conrad et al. (2018). As illustrated in Fig. 1, high-level semantic patterns related to medical conditions can emerge across multiple modalities. Capturing these patterns in a cross-modal, temporal manner is essential for improving predictive performance. Furthermore, critical temporal patterns in EHR data unfold at different time scales Zhang et al. (2023a); Luo et al. (2020); Ma et al. (2020); Ye et al. (2020), yet existing methods struggle to capture variations across these granularities. For instance, sudden changes in vital signs—such as a sharp drop in oxygen saturation or a rapid heart rate increase—may indicate an acute health crisis. In contrast, longer-term trends, such as persistently high blood pressure or a gradual decline in respiratory function, may signal deteriorating health or an impending critical event. Effectively analyzing patient states across multiple time scales is crucial for comprehensive EHR modeling, yet remains an open challenge in current methodologies.
@@
To address these limitations, we propose the Cross-modal Temporal Pattern Discovery (CTPD) framework, designed to extract meaningful temporal patterns from multimodal EHR data to improve the accuracy of clinical outcome predictions. The core innovation of our approach is a novel temporal pattern discovery module, which identifies corresponding temporal patterns (i.e., temporal prototypes) with meaningful semantics across both modalities throughout the dataset. This approach ensures that the model captures essential temporal semantics relevant to patient outcomes, providing a more comprehensive understanding of the data. To further enhance the quality of the learned temporal patterns, we introduce a Temporal Pattern Noise Contrastive Estimation (TP-NCE) loss for aligning pattern embeddings across modalities, along with auxiliary reconstruction losses to ensure that the patterns retain core information of the data. Moreover, our framework incorporates a transformer Vaswani et al. (2017)-based fusion mechanism to effectively fuse the discovered temporal patterns with timestamp-level representations from both modalities, leading to more accurate predictions. We evaluate CTPD on two critical clinical prediction tasks: 48-hour in-hospital mortality prediction and 24-hour phenotype classification, using the MIMIC-III database. The results demonstrate the effectiveness of our approach, which significantly outperforms existing methods, and suggest a promising direction for improving multimodal EHR analysis for clinical prediction.
@@ 2.1 EHR Time-Series Data Analysis
EHR data is critical for clinical tasks such as disease diagnosis, mortality prediction, and treatment planning (Harutyunyan et al., 2019; Zhang et al., 2023b). However, its high dimensionality and irregular nature pose challenges for traditional predictive models (Rani and Sikka, 2012; Lee et al., 2017). Deep learning models, such as RNNs and LSTMs, are often used to capture temporal dependencies in EHR data (Hayat et al., 2022; Deldari et al., 2023), but they struggle with irregular time intervals due to their reliance on fixed-length sequences (Xie et al., 2021). To address this, some methods update patient representations at each time step using graph neural networks (Zhang et al., 2021), while others employ time-aware embeddings to incorporate temporal information (Qian et al., 2023; Zhang et al., 2023b). Despite these advancements, existing approaches still struggle to model high-level temporal patterns essential for accurate clinical outcome prediction.
@@ 2.2 Prototype-based Pattern Learning
Prototype-based learning identifies representative instances (prototypes) and optimizes their distance from input data in latent space for tasks like classification (Li et al., 2023a; Ye et al., 2024). This approach has been widely applied in tasks such as anomaly detection, unsupervised learning, and few-shot learning (Tanwisuth et al., 2021; Li et al., 2023b). Recently, it has been extended to time-series data (Ghosal and Abbasi-Asl, 2021; Li et al., 2023a; Yu et al., 2024), demonstrating its potential for detecting complex temporal patterns. Additionally, prototype-based learning offers interpretable predictions, which is essential for healthcare applications (Zhang et al., 2024). However, learning efficient cross-modal temporal prototypes for multimodal EHR data remains an unexplored problem, as irregular time series and multi-scale patterns present significant challenges for existing methods.
@@ 2.3 Multi-modal Learning in Healthcare
In healthcare, patient data is typically collected in various forms—such as vital signs, laboratory results, medications, medical images, and clinical notes—to provide a comprehensive view of a patient’s health. Integrating these diverse modalities significantly enhances the performance of clinical tasks (Hayat et al., 2022; Zhang et al., 2023b; Yao et al., 2024). However, fusing multimodal data remains challenging due to the heterogeneity and complexity of the sources. Earlier research on multimodal learning (Trong et al., 2020; Ding et al., 2022; Hayat et al., 2022) often rely on late fusion strategies, where unimodal representations are combined via concatenation or Kronecker products. While straightforward, these approaches often fail to capture complex inter-modal interactions, leading to suboptimal representations. Recent works have introduced transformer-based models that focus on cross-modal token interactions (Zhang et al., 2023b; Theodorou et al., 2024; Yao et al., 2024). While these models are effective at capturing inter-modal relationships, they often struggle to extract high-level temporal semantics from multimodal data, limiting the ability to achieve a comprehensive understanding of a patient’s health conditions.
@@ 3.2 Encoding MITS and Clinical Notes
For MITS, we utilize a gating mechanism that dynamically integrates both irregular time series embeddings and imputed regular time series embeddings , following the approach in Zhang et al. (2023b). Formally, the MITS embedding is computed as:
@@
The regular time series embedding is derived by applying a 1D convolution layer to the imputed time series. At each reference time point , the imputed values are sourced from the nearest preceding values or replaced with a standard normal value if no prior data is available. Concurrently, mTAND (multi-time attention) Shukla and Marlin (2021) generates an alternative set of time series representations with the same reference time points with irregular time representations. Specifically, we leverage different Time2Vec Kazemi et al. (2019) functions to produce interpolation embeddings at each time point , which are then concatenated and linearly projected to form .
@@
This yields temporally aligned text embeddings at each time point , enabling effective interaction with time series data. Note that our choice of BERT-Tiny is based on prior work Park et al. (2022) in multimodal EHR analysis, where its lightweight architecture has proven effective. Nonetheless, our framework can be seamlessly integrated with other text encoders.
@@ 3.3 Discover Cross-modal Temporal Patterns from Multimodal EHR
High-level temporal patterns in multimodal EHR data often encode rich medical condition-related semantics that are crucial for predicting clinical outcomes. However, previous works primarily focus on timestamp-level embeddings, frequently overlooking these important temporal patterns (Bahadori and Lipton, 2019; Xiao et al., 2023; Sun et al., 2024). Drawing inspiration from object-centric learning in the computer vision domain (Locatello et al., 2020; Li et al., 2021), we propose a novel temporal pattern discovery module to capture complex patterns within longitudinal data.
Considering the hierarchical nature Yue et al. (2022); Cai et al. (2024) of time series data, the critical temporal patterns for EHR may manifest across multiple time scales. Consequently, our approach performs temporal pattern discovery on multi-scale time series embeddings.
Extracting Cross-modal Temporal Patterns. Owing to the correspondence within multi-modal data, our cross-modal temporal pattern discovery module focuses on extracting corresponding temporal patterns across both modalities for a better understanding of multimodal EHR. Starting with the time series embeddings in Eq. 1, we generate multi-scale embeddings (each scale can be divided by T) using three convolutional blocks followed by mean pooling with a stride of 2 along the time dimension. Each resulting embedding corresponds to a different temporal scale, where the sequence lengths are fractions of the original length (e.g., , , and ). The concatenated multi-scale embedding serves as the diverse temporal representation, where . Since is set to either 48 or 24 in our experiments, remains an integer. We then enhance these embeddings by applying position encoding: , , where denotes the position embeddings in (Vaswani et al., 2017). Furthermore, to capture potential temporal patterns, we define a group of learnable vectors as temporal prototypes, , initially sampled from a normal distribution and refined during training. The shared prototype embeddings are designed to capture semantic-corresponding temporal patterns across modalities, respectively, with and randomly initialized and subsequently optimized.
@@
where and are the prototype embeddings from the previous step , is Gated Recurrent Unit (Cho et al., 2014), and denotes MLP. The above process is repeated for 3 iterations per step. Those refined embeddings denote the discovered temporal patterns for each modality. Note that our design of attention and GRU is inspired by Slot Attention Locatello et al. (2020), which has been shown to be effective in learning object-centric representations from images.
TP-NCE Contraint. To ensure consistent semantics across modalities, we introduce a Temporal Pattern Noise Contrastive Estimation (TP-NCE) loss, inspired by InfoNCE Oord et al. (2018), to enforce the similarity of multimodal prototype embeddings for the same ICU stay while increasing the distance between prototype embeddings from different ICU stays. For a minibatch of samples, the TP-NCE loss from MITS to notes is defined as:
@@
Auxiliary Reconstruction. To ensure that the learned prototype representations capture core information from multimodal EHR data, we introduce two reconstruction objectives aimed at reconstructing imputed regular time series and text embeddings from the learned prototypes. Specifically, we implement a time series decoder to reconstruct the imputed regular time series from , and a text embedding decoder to reconstruct text embeddings from . Both decoders are based on a transformer decoder architecture Vaswani et al. (2017), and two mean squared error (MSE) losses denoted by and are used as the objective function. Here, we define the overall reconstruction loss .
@@ 3.4 Multimodal Fusion
Since information from both modalities is crucial for predicting medical conditions, we propose a multimodal fusion mechanism to integrate these inputs. First, we apply a 2-layer transformer encoder Vaswani et al. (2017) to capture interactions between timestamp-level and prototype embeddings across both modalities for each sample. We continue to use and to represent the resulted prototype embeddings for time-series data and clinical notes, respectively. Similarly, and denote the corresponding timestamp-level embeddings. Then, we aggregate prototype embeddings and timestamp-level embeddings of each modality using an attention-based pooling mechanism:
@@ 4.1 Experimental Setup
Dataset. We assess our model’s efficacy using MIMIC-III v1.4111https://physionet.org/content/mimiciii/1.4/, a comprehensive open-source multimodal clinical database (Johnson et al., 2016). We focus our evaluation on two critical tasks, 48-hour in-hospital mortality prediction (48-IHM) and 24-hour phenotype classification (24-PHE), as established in prior research (Zhang et al., 2023b; Hayat et al., 2022). We extracted raw, irregular time-series data (containing 17 clinical variables) from the MIMIC-III database and selected time series within the first 48-hour and 24-hour windows within each ICU stay for each respective task, as done in Zhang et al. (2023b); Hayat et al. (2022); Harutyunyan et al. (2019). ICU stays shorter than 48 or 24 hours were excluded from our dataset, which explains the different sample size between these two tasks. Unlike Harutyunyan et al. (2019), we did not apply imputation during preprocessing but instead retained the original irregular time-series structure, following Zhang et al. (2023b). Following the dataset splitting by (Harutyunyan et al., 2019), we ensure that the model evaluation is robust by partitioning the data into training, validation, and testing sets, based on unique subject IDs to prevent information leakage. For multimodal analysis, we paired numerical time-series data with corresponding clinical notes from each patient’s ICU stay, consistent with Zhang et al. (2023b). Note that the pipeline of processing clinical notes we used follows the practice in Khadanga et al. (2019). The dataset statistics, including sample counts before and after multimodal paring, are presented in Table 1. The additional details of experimental setup can be found in Appendix A.
@@
| CNN (LeCun et al., 1998) | 85.80 | 49.73 | 46.37 | 75.36 | 38.10 | 40.80 |
| RNN (Elman, 1990) | 84.75 | 46.57 | 45.60 | 73.78 | 36.76 | 33.99 |
| LSTM Hochreiter (1997) | 85.22 | 46.93 | 45.72 | 74.46 | 36.80 | 39.45 |
| Transformer (Vaswani et al., 2017) | 83.45 | 43.03 | 39.31 | 74.98 | 39.37 | 36.13 |
| IP-Net (Shukla and Marlin, 2019) | 81.76 | 39.50 | 43.89 | 73.98 | 35.31 | 39.38 |
| GRU-D (Che et al., 2018) | 49.21 | 12.85 | 19.63 | 52.11 | 17.99 | 26.17 |
| DGM-O (Wu et al., 2021) | 71.99 | 28.67 | 31.72 | 60.70 | 22.56 | 28.87 |
| mTAND (Shukla and Marlin, 2021) | 85.27 | 49.82 | 48.02 | 72.79 | 35.95 | 32.42 |
| SeFT (Horn et al., 2020) | 65.00 | 22.93 | 19.70 | 60.50 | 23.57 | 21.26 |
| UTDE (Zhang et al., 2023b) | 86.14 | 50.60 | 49.29 | 73.62 | 36.80 | 40.58 |
@@
| Flat Deznabi et al. (2021) | 85.71 | 50.96 | 45.80 | 81.77 | 53.79 | 52.13 |
| HierTrans Pappagari et al. (2019) | 84.32 | 47.92 | 42.01 | 80.79 | 51.97 | 50.85 |
| T-LSTM Baytas et al. (2017) | 85.70 | 45.29 | 42.84 | 81.15 | 50.23 | 49.77 |
| FT-LSTM Zhang et al. (2020a) | 84.28 | 43.93 | 36.87 | 81.66 | 51.71 | 50.21 |
| GRU-D (Che et al., 2018) | 72.58 | 26.38 | 31.64 | 49.52 | 16.90 | 27.80 |
| mTAND (Shukla and Marlin, 2021) | 85.40 | 49.68 | 35.76 | 82.14 | 54.57 | 52.01 |
@@
| MMTM (Joze et al., 2020) | 87.88 | 53.58 | 51.54 | 81.46 | 51.88 | 51.59 |
| DAFT (Pölsterl et al., 2021) | 87.53 | 52.40 | 51.95 | 81.18 | 50.91 | 50.72 |
| MedFuse (Hayat et al., 2022) | 86.02 | 51.00 | 49.29 | 78.88 | 45.99 | 47.47 |
| DrFuse (Yao et al., 2024) | 85.97 | 49.94 | 49.75 | 80.88 | 49.62 | 50.18 |

### v3

@@ Abstract
Integrating multimodal Electronic Health Records (EHR) data—such as numerical time series and free-text clinical reports—has great potential in predicting clinical outcomes. However, prior work has primarily focused on capturing temporal interactions within individual samples and fusing multimodal information, overlooking critical temporal patterns across patients. These patterns, such as trends in vital signs like abnormal heart rate or blood pressure, can indicate deteriorating health or an impending critical event. Similarly, clinical notes often contain textual descriptions that reflect these patterns. Identifying corresponding temporal patterns across different modalities is crucial for improving the accuracy of clinical outcome predictions, yet it remains a challenging task. To address this gap, we introduce a Cross-Modal Temporal Pattern Discovery (CTPD) framework, designed to efficiently extract meaningful cross-modal temporal patterns from multimodal EHR data. Our approach introduces shared initial temporal pattern representations which are refined using slot attention to generate temporal semantic embeddings. To ensure rich cross-modal temporal semantics in the learned patterns, we introduce a Temporal Pattern Noise Contrastive Estimation (TP-NCE) loss for cross-modal alignment, along with two reconstruction losses to retain core information of each modality. Evaluations on two clinically critical tasks—48-hour in-hospital mortality and 24-hour phenotype classification—using the MIMIC-III database demonstrate the superiority of our method over existing approaches. The code is available at https://github.com/HKU-MedAI/CTPD.
@@
Fuying Wang1∗, Feng Wu1††thanks: Equal contributions., Yihan Tang1, Lequan Yu1††thanks: Corresponding Author. 1Department of Statistics and Actuarial Science, School of Computing and Data Science, The University of Hong Kong {fuyingw,fengwu96,hank-tang}@connect.hku.hk, lqyu@hku.hk Correspondence: lqyu@hku.hk
@@ 1 Introduction
The increasing availability of Electronic Health Records (EHR) presents significant opportunities for advancing predictive modeling in healthcare (acosta2022; wang2024multimodal). EHR data is inherently multimodal and time-aware, encompassing structured data like vital signs, laboratory results, and medications, as well as unstructured data such as free-text clinical reports kim2023heterogeneous. Integrating these diverse data types is crucial for comprehensive patient monitoring and accurate prediction of clinical outcomes hayat2022medfuse; wang2023hierarchical; zhang2023improving. However, the irregularity and heterogeneity of multimodal data present significant challenges for precise outcome prediction.
@@
Existing approaches primarily address either the irregularity of data shukla2019interpolation; horn2020set; zhang2021graph; zhang2023improving or the fusion of multiple modalities huang2020multimodal; zhang2020advances; xu2021mufasa; kline2022, but they often neglect broader temporal trends that span across patient cases. These cross-modal temporal patterns, present in both structured and unstructured data, can provide high-level semantic insights into a patient’s health trajectory and potential risks conrad2018temporal. As illustrated in Fig. 1, high-level semantic patterns related to medical conditions can emerge across multiple modalities. Capturing these patterns in a cross-modal, temporal manner is essential for improving predictive performance. Furthermore, critical temporal patterns in EHR data unfold at different time scales zhang2023warpformer; luo2020hitanet; ma2020adacare; ye2020lsan, yet existing methods struggle to capture variations across these granularities. For instance, sudden changes in vital signs—such as a sharp drop in oxygen saturation or a rapid heart rate increase—may indicate an acute health crisis. In contrast, longer-term trends, such as persistently high blood pressure or a gradual decline in respiratory function, may signal deteriorating health or an impending critical event. Effectively analyzing patient states across multiple time scales is crucial for comprehensive EHR modeling, yet remains an open challenge in current methodologies.
@@
To address these limitations, we propose the Cross-modal Temporal Pattern Discovery (CTPD) framework, designed to extract meaningful temporal patterns from multimodal EHR data to improve the accuracy of clinical outcome predictions. The core innovation of our approach is a novel temporal pattern discovery module, which identifies corresponding temporal patterns (i.e., temporal prototypes) with meaningful semantics across both modalities throughout the dataset. This approach ensures that the model captures essential temporal semantics relevant to patient outcomes, providing a more comprehensive understanding of the data. To further enhance the quality of the learned temporal patterns, we introduce a Temporal Pattern Noise Contrastive Estimation (TP-NCE) loss for aligning pattern embeddings across modalities, along with auxiliary reconstruction losses to ensure that the patterns retain core information of the data. Moreover, our framework incorporates a transformer vaswani2017attention-based fusion mechanism to effectively fuse the discovered temporal patterns with timestamp-level representations from both modalities, leading to more accurate predictions. We evaluate CTPD on two critical clinical prediction tasks: 48-hour in-hospital mortality prediction and 24-hour phenotype classification, using the MIMIC-III database. The results demonstrate the effectiveness of our approach, which significantly outperforms existing methods, and suggest a promising direction for improving multimodal EHR analysis for clinical prediction.
@@ 2.1 EHR Time-Series Data Analysis
EHR data is critical for clinical tasks such as disease diagnosis, mortality prediction, and treatment planning (harutyunyan2019multitask; zhang2023improving). However, its high dimensionality and irregular nature pose challenges for traditional predictive models (rani2012recent; lee2017big). Deep learning models, such as RNNs and LSTMs, are often used to capture temporal dependencies in EHR data (hayat2022medfuse; deldari2023latent), but they struggle with irregular time intervals due to their reliance on fixed-length sequences (xie2021). To address this, some methods update patient representations at each time step using graph neural networks (zhang2021graph), while others employ time-aware embeddings to incorporate temporal information (qian2023knowledge; zhang2023improving). Despite these advancements, existing approaches still struggle to model high-level temporal patterns essential for accurate clinical outcome prediction.
@@ 2.2 Prototype-based Pattern Learning
Prototype-based learning identifies representative instances (prototypes) and optimizes their distance from input data in latent space for tasks like classification (li2023prototypes; ye2024ptarl). This approach has been widely applied in tasks such as anomaly detection, unsupervised learning, and few-shot learning (tanwisuth2021prototype; li2023prototype). Recently, it has been extended to time-series data (ghosal2021multi; li2023prototypes; yu2024imputation), demonstrating its potential for detecting complex temporal patterns. Additionally, prototype-based learning offers interpretable predictions, which is essential for healthcare applications (zhang2024prototypical). However, learning efficient cross-modal temporal prototypes for multimodal EHR data remains an unexplored problem, as irregular time series and multi-scale patterns present significant challenges for existing methods.
@@ 2.3 Multi-modal Learning in Healthcare
In healthcare, patient data is typically collected in various forms—such as vital signs, laboratory results, medications, medical images, and clinical notes—to provide a comprehensive view of a patient’s health. Integrating these diverse modalities significantly enhances the performance of clinical tasks (hayat2022medfuse; zhang2023improving; yao2024drfuse). However, fusing multimodal data remains challenging due to the heterogeneity and complexity of the sources. Earlier research on multimodal learning (trong2020late; ding2022multimodal; hayat2022medfuse) often rely on late fusion strategies, where unimodal representations are combined via concatenation or Kronecker products. While straightforward, these approaches often fail to capture complex inter-modal interactions, leading to suboptimal representations. Recent works have introduced transformer-based models that focus on cross-modal token interactions (zhang2023improving; theodorou2024framm; yao2024drfuse). While these models are effective at capturing inter-modal relationships, they often struggle to extract high-level temporal semantics from multimodal data, limiting the ability to achieve a comprehensive understanding of a patient’s health conditions.
@@ 3.2 Encoding MITS and Clinical Notes
For MITS, we utilize a gating mechanism that dynamically integrates both irregular time series embeddings and imputed regular time series embeddings , following the approach in zhang2023improving. Formally, the MITS embedding is computed as:
@@
The regular time series embedding is derived by applying a 1D convolution layer to the imputed time series. At each reference time point , the imputed values are sourced from the nearest preceding values or replaced with a standard normal value if no prior data is available. Concurrently, mTAND (multi-time attention) shukla2021multi generates an alternative set of time series representations with the same reference time points with irregular time representations. Specifically, we leverage different Time2Vec kazemi2019time2vec functions to produce interpolation embeddings at each time point , which are then concatenated and linearly projected to form .
@@
This yields temporally aligned text embeddings at each time point , enabling effective interaction with time series data. Note that our choice of BERT-Tiny is based on prior work park2022graph in multimodal EHR analysis, where its lightweight architecture has proven effective. Nonetheless, our framework can be seamlessly integrated with other text encoders.
@@ 3.3 Discover Cross-modal Temporal Patterns from Multimodal EHR
High-level temporal patterns in multimodal EHR data often encode rich medical condition-related semantics that are crucial for predicting clinical outcomes. However, previous works primarily focus on timestamp-level embeddings, frequently overlooking these important temporal patterns (bahadori2019temporal; xiao2023gaformer; sun2024time). Drawing inspiration from object-centric learning in the computer vision domain (locatello2020object; li2021scouter), we propose a novel temporal pattern discovery module to capture complex patterns within longitudinal data.
Considering the hierarchical nature yue2022ts2vec; cai2024msgnet of time series data, the critical temporal patterns for EHR may manifest across multiple time scales. Consequently, our approach performs temporal pattern discovery on multi-scale time series embeddings.
Extracting Cross-modal Temporal Patterns. Owing to the correspondence within multi-modal data, our cross-modal temporal pattern discovery module focuses on extracting corresponding temporal patterns across both modalities for a better understanding of multimodal EHR. Starting with the time series embeddings in Eq. 1, we generate multi-scale embeddings (each scale can be divided by T) using three convolutional blocks followed by mean pooling with a stride of 2 along the time dimension. Each resulting embedding corresponds to a different temporal scale, where the sequence lengths are fractions of the original length (e.g., , , and ). The concatenated multi-scale embedding serves as the diverse temporal representation, where . Since is set to either 48 or 24 in our experiments, remains an integer. We then enhance these embeddings by applying position encoding: , , where denotes the position embeddings in (vaswani2017attention). Furthermore, to capture potential temporal patterns, we define a group of learnable vectors as temporal prototypes, , initially sampled from a normal distribution and refined during training. The shared prototype embeddings are designed to capture semantic-corresponding temporal patterns across modalities, respectively, with and randomly initialized and subsequently optimized.
@@
where and are the prototype embeddings from the previous step , is Gated Recurrent Unit (cho2014learning), and denotes MLP. The above process is repeated for 3 iterations per step. Those refined embeddings denote the discovered temporal patterns for each modality. Note that our design of attention and GRU is inspired by Slot Attention locatello2020object, which has been shown to be effective in learning object-centric representations from images.
TP-NCE Contraint. To ensure consistent semantics across modalities, we introduce a Temporal Pattern Noise Contrastive Estimation (TP-NCE) loss, inspired by InfoNCE oord2018representation, to enforce the similarity of multimodal prototype embeddings for the same ICU stay while increasing the distance between prototype embeddings from different ICU stays. For a minibatch of samples, the TP-NCE loss from MITS to notes is defined as:
@@
Auxiliary Reconstruction. To ensure that the learned prototype representations capture core information from multimodal EHR data, we introduce two reconstruction objectives aimed at reconstructing imputed regular time series and text embeddings from the learned prototypes. Specifically, we implement a time series decoder to reconstruct the imputed regular time series from , and a text embedding decoder to reconstruct text embeddings from . Both decoders are based on a transformer decoder architecture vaswani2017attention, and two mean squared error (MSE) losses denoted by and are used as the objective function. Here, we define the overall reconstruction loss .
@@ 3.4 Multimodal Fusion
Since information from both modalities is crucial for predicting medical conditions, we propose a multimodal fusion mechanism to integrate these inputs. First, we apply a 2-layer transformer encoder vaswani2017attention to capture interactions between timestamp-level and prototype embeddings across both modalities for each sample. We continue to use and to represent the resulted prototype embeddings for time-series data and clinical notes, respectively. Similarly, and denote the corresponding timestamp-level embeddings. Then, we aggregate prototype embeddings and timestamp-level embeddings of each modality using an attention-based pooling mechanism:
@@ 4.1 Experimental Setup
Dataset. We assess our model’s efficacy using MIMIC-III v1.4111https://physionet.org/content/mimiciii/1.4/, a comprehensive open-source multimodal clinical database (johnson2016mimic). We focus our evaluation on two critical tasks, 48-hour in-hospital mortality prediction (48-IHM) and 24-hour phenotype classification (24-PHE), as established in prior research (zhang2023improving; hayat2022medfuse). We extracted raw, irregular time-series data (containing 17 clinical variables) from the MIMIC-III database and selected time series within the first 48-hour and 24-hour windows within each ICU stay for each respective task, as done in zhang2023improving; hayat2022medfuse; harutyunyan2019multitask. ICU stays shorter than 48 or 24 hours were excluded from our dataset, which explains the different sample size between these two tasks. Unlike harutyunyan2019multitask, we did not apply imputation during preprocessing but instead retained the original irregular time-series structure, following zhang2023improving. Following the dataset splitting by (harutyunyan2019multitask), we ensure that the model evaluation is robust by partitioning the data into training, validation, and testing sets, based on unique subject IDs to prevent information leakage. For multimodal analysis, we paired numerical time-series data with corresponding clinical notes from each patient’s ICU stay, consistent with zhang2023improving. Note that the pipeline of processing clinical notes we used follows the practice in khadanga2019using. The dataset statistics, including sample counts before and after multimodal paring, are presented in Table 1. The additional details of experimental setup can be found in Appendix A.
@@
| CNN (lecun1998gradient) | 85.80 | 49.73 | 46.37 | 75.36 | 38.10 | 40.80 |
| RNN (elman1990finding) | 84.75 | 46.57 | 45.60 | 73.78 | 36.76 | 33.99 |
| LSTM hochreiter1997long | 85.22 | 46.93 | 45.72 | 74.46 | 36.80 | 39.45 |
| Transformer (vaswani2017attention) | 83.45 | 43.03 | 39.31 | 74.98 | 39.37 | 36.13 |
| IP-Net (shukla2019interpolation) | 81.76 | 39.50 | 43.89 | 73.98 | 35.31 | 39.38 |
| GRU-D (che2018recurrent) | 49.21 | 12.85 | 19.63 | 52.11 | 17.99 | 26.17 |
| DGM-O (wu2021dynamic) | 71.99 | 28.67 | 31.72 | 60.70 | 22.56 | 28.87 |
| mTAND (shukla2021multi) | 85.27 | 49.82 | 48.02 | 72.79 | 35.95 | 32.42 |
| SeFT (horn2020set) | 65.00 | 22.93 | 19.70 | 60.50 | 23.57 | 21.26 |
| UTDE (zhang2023improving) | 86.14 | 50.60 | 49.29 | 73.62 | 36.80 | 40.58 |
@@
| Flat deznabi2021predicting | 85.71 | 50.96 | 45.80 | 81.77 | 53.79 | 52.13 |
| HierTrans pappagari2019hierarchical | 84.32 | 47.92 | 42.01 | 80.79 | 51.97 | 50.85 |
| T-LSTM baytas2017patient | 85.70 | 45.29 | 42.84 | 81.15 | 50.23 | 49.77 |
| FT-LSTM zhang2020time | 84.28 | 43.93 | 36.87 | 81.66 | 51.71 | 50.21 |
| GRU-D (che2018recurrent) | 72.58 | 26.38 | 31.64 | 49.52 | 16.90 | 27.80 |
| mTAND (shukla2021multi) | 85.40 | 49.68 | 35.76 | 82.14 | 54.57 | 52.01 |
@@
| MMTM (joze2020mmtm) | 87.88 | 53.58 | 51.54 | 81.46 | 51.88 | 51.59 |
| DAFT (polsterl2021combining) | 87.53 | 52.40 | 51.95 | 81.18 | 50.91 | 50.72 |
| MedFuse (hayat2022medfuse) | 86.02 | 51.00 | 49.29 | 78.88 | 45.99 | 47.47 |
| DrFuse (yao2024drfuse) | 85.97 | 49.94 | 49.75 | 80.88 | 49.62 | 50.18 |